# CSGO: Content-Style Composition in Text-to-Image Generation

### v1

@@ 4.1 CSGO framework
where $\varepsilon$ denotes the random sampled Gaussian noise, $\varepsilon_{\theta}$ denotes the trainable parameters of CSGO, $t$ represents the timestep. Note that the latent latent $z_{t}$ is constructed with a styl image $T$ during training, $z_{t}=\sqrt{\bar{\alpha}_{t}}\psi(T)+\sqrt{1-\bar{\alpha}_{t}}\varepsilon$, where $\psi(\cdot)$ mapping the original input to the latent space function, $\bar{\alpha_{t}}$ is consistent with diffusion models Song et al. (2020); Ho et al. (2020). We randomly drop content image and style image conditions in the training phase to enable classifier-free guidance in the inference stage.
@@ 5.2 Experimental Results
In terms of content retention, it can be observed that StyleID Chung et al. (2024) and StyleAligned Hertz et al. (2024), which are based on inversion, maintain the original content too strongly in sketch style transfer scenarios (CAS is very low). However. they are unable to inject style information since CSD score is low. InstantStyle Wang et al. (2024a) and StyleShot Junyao et al. (2024) (including Lineart), which use lines to control the content, are affected by the level of detail of the lines, and have different degrees of loss of content (such as face scenes). The proposed CSGO directly utilizes all the information of the content image, and content preservation is optimal. The quantitative results in Table 1 also show that the proposed CSGO maintains high-quality content retention with precise style transfer.
@@
Text-Driven Stylized Synthesis. The proposed method enables text-driven style control, i.e., given a text prompt and a style image, generating images with similar styles. Figure 6 shows the comparison of the generation results of the proposed CSGO with the state-of-the-art methods. In a simple scene, it is intuitive to observe that our CSGO obeys textual prompts more. The reason for this is that thanks to the explicit decoupling of content and style features, style images only inject style information without exposing content. In addition, in the complex scene, thanks to the well-designed style feature injection block, CSGO enables optimal style control while converting the meaning of text. As illustrated in Figure 8, we demonstrated more results.
@@
Content control and style control. We discuss the impact of the two feature injection methods, as shown in Figure 10. The content image must be injected via ControlNet injection to maintain the layout while preserving the semantic information. Content features only retain semantic information if injected into the CSGO framework only through IP-Adapter injection (Figure 10(1)).
@@
After introducing the ControlNet injection, the quality of content retention improved, as shown in Figure 12. However, if the style features are injected into base UNet only without ControlNet injection, this weakens the style of the generated images, which can be observed in the comparison of Figure 10(2) and (3). Therefore, the proposed CSGO pre-injects style features in the ControlNet branch to further fuse the style features to enhance the transfer effect.
@@
Token number. We explore the effect of the number of token $t$ in the style projection layer on the results of style transfer and text-driven style synthesis. The experimental results are shown in Figure 10, where the style control becomes progressively better as $t$ increases. This is in line with our expectation that $t$ influences the quality of feature extraction. A larger $t$ means that the projection layer is able to extract richer style features.

### v2

@@ 4.1 CSGO framework
where $\varepsilon$ denotes the random sampled Gaussian noise, $\varepsilon_{\theta}$ denotes the trainable parameters of CSGO, $t$ represents the timestep. Note that the latent latent $z_{t}$ is constructed with a style image $T$ during training, $z_{t}=\sqrt{\bar{\alpha}_{t}}\psi(T)+\sqrt{1-\bar{\alpha}_{t}}\varepsilon$, where $\psi(\cdot)$ mapping the original input to the latent space function, $\bar{\alpha_{t}}$ is consistent with diffusion models Song et al. (2020); Ho et al. (2020). We randomly drop content image and style image conditions in the training phase to enable classifier-free guidance in the inference stage.
@@ 5.2 Experimental Results
In terms of content retention, it can be observed that StyleID Chung et al. (2024) and StyleAligned Hertz et al. (2024), which are based on inversion, maintain the original content too strongly in sketch style transfer scenarios (CAS is very low). However. they are unable to inject style information since CSD score is low. InstantStyle Wang et al. (2024a) and StyleShot Junyao et al. (2024) (including Lineart), which use lines to control the content, are affected by the level of detail of the lines and have different degrees of loss of content (such as face scenes). The proposed CSGO directly utilizes all the information of the content image, and content preservation is optimal. The quantitative results in Table 1 also show that the proposed CSGO maintains high-quality content retention with precise style transfer.
Text-Driven Stylized Synthesis. The proposed method enables text-driven style control, i.e., given a text prompt and a style image, generating images with similar styles. Figure 6 shows the comparison of the generation results of the proposed CSGO with the state-of-the-art methods. In a simple scene, it is intuitive to observe that our CSGO obeys textual prompts more. The reason for this is that thanks to the explicit decoupling of content and style features, style images only inject style information without exposing content. In addition, in complex scenes, thanks to the well-designed style feature injection block, CSGO enables optimal style control while converting the meaning of text. As illustrated in Figure 8, we demonstrated more results.
@@
Content control and style control. We discuss the impact of the two feature injection methods, as shown in Figure 10. The content image must be injected via ControlNet injection to maintain the layout while preserving the semantic information. If content images are injected into the base model only through an additional Cross attention layer, only semantic information is guaranteed, while the full content information is not preserved (Figure 10(1)). After introducing the ControlNet injection, the quality of content retention improved, as shown in Figure 12. However, if the style features are injected into base UNet only without ControlNet injection, this weakens the style of the generated images, which can be observed in the comparison of Figure 10(2) and (3). Therefore, the proposed CSGO pre-injects style features in the ControlNet branch to further fuse the style features to enhance the transfer effect.
@@
Token number. We explore the effect of the number of token $t$ in the style projection layer on the results of style transfer and text-driven style synthesis. The experimental results are shown in Figure 10, where the style control becomes progressively better as $t$ increases. This is in line with our expectation that $t$ influences the quality of feature extraction. A larger $t$ means that the projection layer can extract richer style features.